OpenStudy (mcba1234):

I have to write this expression as a single natural logarithm. 3 ln 3 + 3 ln c I got ln27(c^3) is this right?

OpenStudy (s4sensitiveandshy):

\[ a~ \log ~b = \log b^a \]\[\log x + \log y = \log (x * y)\]

OpenStudy (s4sensitiveandshy):

good job.